OpenStudy (anonymous):

What are the rules of exponents within the Brackets and and on the outside of the brackets

OpenStudy (anonymous):

(x * y)^2 = x^2 * y^2 x*y^2 = x * y^2 This might help. Not sure exactly what you're looking for

OpenStudy (anonymous):

Example Whats the difference between the exponent inside or out side \[(4^4) (4)^4\]

OpenStudy (anonymous):

There is no difference if there is just one thing inside the brackets. Both simplify to just 4^4.

OpenStudy (anonymous):

If there are multiple things multiplied inside the brackets, you have to do it to each one of them. If there are multiple things added inside the brackets, you have to use the distributive property. There are many rules.